OpenStudy (anonymous):

if a(1) =-6 and a(2) =30 what is geometric sequence

OpenStudy (anonymous):

\[a_{n}=(-6)(-5)^{n-1}\]

OpenStudy (anonymous):

\(a_1,a_2,a_3,a_4,...\) since it is geometric it must be \(a_1, a_1r,a_1r^2,a_1r^3,...\) you know \(a_1=-6\) and \(r=\frac{a_2}{a_1}=\frac{30}{-5}=-6\)

OpenStudy (anonymous):

@satellite73 it is -5